# FLARE: Towards Universal Dataset Purification against Backdoor Attacks

### v1

@@ III-B Revisiting Strategy 1: Early Convergence
III-C Revisiting Strategy 2: Dominant Trigger Effects
Researchers in [6, 16] observed that poisoned samples often exhibit highly localized and small saliency regions, suggesting that trigger-related features can be regarded as ‘short-cut’ that are more easily learned by DNNs compared to benign features.
Settings. We use BadNets and WaNet for our discussions, following the same experimental settings as described in Section III-B, with WaNet results limited to the A2A setting for brevity. We compare saliency regions from two perspectives: visualization and quantitative measurement. Specifically, we utilize Grad-CAM [41] to visualize the saliency regions of both benign and poisoned samples. We then calculate the $\ell^{1}$ norms of these regions to quantify their size.
@@ III-C Revisiting Strategy 2: Dominant Trigger Effects
IBD-PSC            SCALE-UP
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
@@
Figure 4: Difference in prediction confidences for benign and poisoned samples on CIFAR-10 under input-level and weight-level perturbations, calculated by comparing prediction confidences on the original label before and after perturbation.
@@
Figure 5: T-SNE visualization of latent representations across different hidden layers on the CIFAR-10 dataset. Each point corresponds to a training sample with poisoned samples marked in red and benign samples in blue.
III-D Revisiting Strategy 3: Perturbation Consistency
Researchers also observed that poisoned samples exhibit greater prediction consistency than those of benign samples under pixel-level amplification [12, 37] or weight-level alterations [39, 15, 52]. This implies that DNNs tend to overfit on triggers instead of learning semantic features.
@@ III-D Revisiting Strategy 3: Perturbation Consistency
Results. As shown in Figure 4, for BadNets (A2O), the differences are consistently close to zero under the pixel-level and weight-level perturbations. It indicates that the predictions of these poisoned samples are unaffected by such perturbations. In contrast, under A2A and UT attacks, the confidence differences of poisoned samples approach 1, indicating that the perturbations significantly change the predictions of poisoned samples. The sensitivity of poisoned samples to perturbations, similar to that of benign samples, indicates that DNNs struggle to overfit to poisoned samples in A2A and UT attacks. Thus, the assumption that backdoor connections are easier to learn than benign ones does not hold for A2A and UT attacks.
@@ III-E Revisiting Latent Separability on a Particular Layer
Previous studies [4, 31] explored the latent separability between benign and poisoned samples in the feature space. In particular, these works primarily focused on the final hidden layer as a representative intermediate layer. We argue that their success also partly relied on the assumption that backdoor can be easily learned. In this section, we verify it.
Settings. We also conduct experiments using BadNets and WaNet on the CIFAR-10 dataset to facilitate our analysis. To assess latent separability, we employ t-SNE visualization to examine the latent representations of both benign and poisoned samples across various hidden layers, with particular focus on shallow, middle, and deep layers. All other experimental settings remain consistent with those outlined in Section III-C.
@@ IV-B Latent Representation Extraction
Figure 6: The main pipeline of FLARE. Stage 1: Latent Representation Extraction: A backdoored model is trained on a poisoned dataset, and each training sample $x_{i}$ is forwarded to generate feature maps at hidden layers. The value ranges of each feature map are aligned using the statistics of corresponding BN layers. An abnormal value is then extracted from each aligned feature map, and these values are aggregated across all hidden layers to form the final representation $\mathcal{R}_{i}$. Stage 2: poisoned samples Detection: With the representations $\mathcal{R}$ of all training samples, UMAP reduces $\mathcal{R}$ to $\hat{\mathcal{R}}$, followed by density-based clustering to separate samples into two clusters. The cluster with a higher $\kappa^{C}$ is identified as poisoned.
@@ IV-B2 Extracting Abnormal Features
As backdoor-related features usually function as dominant features, they tend to induce abnormally large or small activations in the feature map, as partly supported by findings in [2]. Building on this understanding, we propose to focus on identifying outliers within each feature map, specifically targeting the abnormally small or large values. In particular, all feature maps are normalized and their values adhere to a uniform distribution defined by the statistics of the BN layers after the previous alignment step. At this time, both abnormally large and small values fall outside the central distribution, corresponding to regions with low occurrence probability. This allows us to consistently extract the minimum values from each aligned feature map as indicators of backdoor-related features, without needing to separately select the largest or smallest values. Specifically, we can extract the minimum value (i.e., $r_{c}^{(l)}$) from the aligned output $\hat{a}^{(l)}$ of the $l$-th hidden layer via:
@@ IV-C1 Cluster Splitting
[…]
@@ IV-C2 Stable Subspace Selection
Given that poisoned samples typically share the same trigger-related features, they tend to aggregate into a more stable cluster. In contrast, benign samples, which originate from a natural and diverse distribution, generally form clusters with lower stability. Therefore, cluster stability forms an effective indicator for identifying poisoned samples. Traditionally, in HDBSCAN, cluster stability is calculated by aggregating the density level $\lambda$ of each sample within a cluster. However, in poisoned sample detection, where the number of poisoned samples is much smaller than that of benign samples, this approach introduces bias toward the majority class (benign samples). To alleviate this problem, we redefine cluster stability to focus on the persistence of a cluster across varying density levels ($\lambda$), as follows.
@@
To determine the optimal number of layers $k$ to exclude, FLARE designs an adaptive algorithm that dynamically selects a suitable $k$. The algorithm begins with a model configured with all $L$ hidden layers and progressively removes the last hidden layer of the modified model at each step. At each iteration, FLARE splits the condensed tree $\mathbf{T}$ at the root node to form two primary clusters and assesses the stability $\kappa^{C_{large}}$ of the larger cluster $C_{large}$. If the stability $\kappa^{C_{large}}$ surpasses a predefined threshold $\xi$, it suggests that the benign samples are cohesively grouped within the current subspace. To further ensure the cluster stability against potential anomalies, FALRE generates a condensed tree at each step and traverses from the root node of $C_{large}$ down to depth $d$. If the stability at any depth exceeds $\xi$, the current subspace is considered stable.
@@ IV-C3 Stability-based Detection
Within this obtained subspace, FLARE splits the condensed tree $\mathbf{T}$ of the entire poisoned dataset at the root node to form two primary clusters, $C_{1}$ and $C_{2}$. FLARE then evaluates the cluster stability $\kappa^{C}$ of each cluster and identifies the cluster with the highest stability as poisoned. The set of poisoned samples $\mathcal{S}_{p}$ is:
@@ V-A Main Settings
Datasets and Models. We conduct experiments on two classical image classification benchmark datasets, including CIFAR-10 [19] and Tiny-ImageNet [7]. For the primary experiments, we employ ResNet-18 [14] as the default architecture. For ablation studies, we use VGG19 [42] and MobileNetV2 [40].
Attacks Configurations. To validate the generalizability of our defense, we evaluate it against four representative types of backdoor attacks across three attack modes. These attacks are categorized as follows: (1) sample-agnostic attacks: BadNet [11], Blend [5], Trojan [29]; (2) sample-specific attacks: IAD [35], WaNet [36], ISSBA [25]; (3) clean-label attack: LC [44]; (4) sparse attack: SIBA [9]. The attack modes include the all-to-one (A2O), all-to-all (A2A), and untargeted (UT) paradigms. To comprehensively assess each attack mode, attacks are adapted accordingly. For example, BadNets is evaluated in three formats: BadNets (A2O), BadNets (A2A), and BadNets (UT). These attacks are implemented using the open-source BackdoorBox toolkit [24], strictly following the settings outlined in their papers. For all attacks, we set the target label $y_{t}$ to 0. Given the complexity of the Tiny-ImageNet dataset, we focus on the most representative attacks, omitting some similar-effect attacks to avoid excessive computational costs. To ensure consistently high attack success rates, we set the poisoning rate to 0.1 and the cover rate to 0.2.
TABLE I: The detection performance (%) of FLARE on the CIFAR-10 dataset using the ResNet-18 model. The best results (highest TPRs and lowest FPRs) are in bold, while TPRs $<$ 80% and FPRs $>$ 20% are marked in red as fail cases.
@@
TABLE II: The detection performance (%) of FLARE on the TinyImageNet dataset using the ResNet-18 model. The best results (highest TPRs and lowest FPRs) are in bold, while TPRs $<$ 80% and FPRs $>$ 20% are marked in red as fail cases.
@@
TABLE III: Effectiveness of secure training from scratch (FLARE (P)) on the “purified” CIFAR-10 dataset. The best results (highest BAs and lowest ASRs) are in bold, while BAs $<$ 80% and ASRs $>$ 20% are marked in red as fail cases.
@@
TABLE IV: Effectiveness of secure training from scratch (FLARE (P)) on the “purified” Tiny-ImageNet dataset. The best results (highest BAs and lowest ASRs) are in bold, while BAs $<$ 50% and ASRs $>$ 20% are marked in red as fail cases.
@@
Baseline Defenses. To evaluate the detection performance, we compare our FLARE method with five state-of-the-art (SOTA) detection methods, including AC [4], SCALE-UP [12], MSPC [37], IBD-PSC [15], and CT [39]. Following the detection, we further compare our defense performance on the purified dataset using FLARE, dubbed “FLARE (P)”, with these five methods. For backdoor mitigation, we compare our backdoor removal defense via FLARE, dubbed “FLARE (R)”,with six SOTA methods, including FP [28], NAD [21], AMW [3], ABL [20], SEAM [56], and BTI-DBF [50]. All defenses are implemented using the official source codes with the default settings.
V-B Detection Performance
We mainly assess the effectiveness of FLARE against various backdoor attacks on the CIFAR-10 and Tiny-ImageNet datasets. The results, detailed in Tables I and II, demonstrate that FLARE consistently achieves promising performance across different attack scenarios, with the TPRs reaching 100% (or near 100%) while maintaining FPRs near 0%. The results demonstrate a significant improvement over baseline defenses, which fail in A2A and UT attack scenarios (highlighted in red). This failure is primarily due to their implicit assumptions that backdoor correlations are easier to learn than benign ones, which does not hold in more complex attack configurations. Additionally, FLARE maintains an FPR close to zero on datasets composed entirely of benign samples, indicating that it has minimal false positives.
V-C Effectiveness of Secure Training from Scratch
To assess the effectiveness of FLARE (P), which retrains models on a purified dataset with detected poisoned samples removed, we evaluated the BAs and ASRs of the retrained models on the CIFAR-10 and Tiny-ImageNet datasets. As shown in Tables III-IV, the ASRs of retrained models approach to nearly 0% in almost all cases. In contrast, all baseline methods fail in most cases, especially under all-to-all and untargeted attacks. In particular, we also observe some outliers with a high ASR, e.g., models under Blend (UT) and SIBA (UT) on CIFAR-10. This is mainly because even adding triggers of these attacks to a backdoor-free benign model may result in some ASR, instead of due to the failure of our method.
V-D Effectiveness of Backdoor Removal
To assess the effectiveness of FLARE (R), which directly removes backdoors by unlearning the detected poisoned samples, we also evaluated the BAs and ASRs of the purified models on the CIFAR-10 and Tiny-ImageNet datasets. As shown in Table V and Table VI, FLARE (R) significantly reduces the ASRs of different attacks to nearly 0%, while the impact on BA remains minimal, generally around 1%. Compared to existing advanced defenses, FLARE (R) consistently proves effective across various attacks and datasets.
V-E Ablation Study
We conduct ablation studies to assess the impact of four specific factors on the effectiveness of FLARE: (1) target label for A2O attacks, (2) model architectures, 3) the module for selecting a stable subspace, 4) the hyper-parameters of the lambda threshold $\xi$ and maximum depth $d$.
Impact of Model Architectures. In our main results, the model architecture is initially set to ResNet-18. To further assess the robustness of FLARE, we also evaluate it with VGG-19 and MobileNet v2 architectures. The results, shown in Table VII, indicate that FLARE consistently achieves high TPRs, approaching 100%, while maintaining FPRs near 0%.
@@ V-E Ablation Study
Figure 7: Performance of our defense across different target labels of CIFAR-10.
(a) BadNets (A2O)
(b) ISSBA (A2O)
(c) LC
(d) BadNets (A2A)
(e) BadNets (UT)
Figure 8: The impact of poisoning rate on CIFAR-10.
TABLE V: Comparisons of FLARE (R) with SOTA backdoor-removal defenses on CIFAR-10. The best results (highest BAs and lowest ASRs) are in bold, while BAs $<$ 80% and ASRs $>$ 20% are marked in red as fail cases.

### v2

@@ III-B Revisiting Strategy 1 (Early Convergence)
III-C Revisiting Strategy 2 (Dominant Trigger Effects)
Researchers in chou2020sentinet ; huang2023distilling observed that poisoned samples often exhibit highly localized and small saliency regions, suggesting that trigger-related features can be regarded as ‘short-cut’ that are more easily learned by DNNs compared to benign features.
Settings. We use BadNets and WaNet for our discussions, following the same experimental settings as described in Section III-B, with WaNet results limited to the A2A setting for brevity. We compare saliency regions from two perspectives: visualization and quantitative measurement. Specifically, we utilize Grad-CAM selvaraju2017grad to visualize the saliency regions of both benign and poisoned samples. We then calculate the $\ell^{1}$ norms of these regions to quantify their size.
@@ III-C Revisiting Strategy 2 (Dominant Trigger Effects)
CT              IBD-PSC              MSPC              SCALE-UP
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 4: Difference in prediction confidences for benign and poisoned samples on CIFAR-10 under input-level and weight-level perturbations, calculated by comparing prediction confidences on the original label before and after perturbation. A significant difference in confidence patterns indicates effective detection, as poisoned samples are distinguishable from benign ones.
@@
Figure 5: T-SNE visualization of latent representations across different hidden layers on the CIFAR-10 dataset. A total of 2,000 samples are randomly sampled from the training set, including both benign and poisoned samples, with a poisoning rate of 0.1. Each point corresponds to a training sample with poisoned samples marked in red and benign samples in blue.
III-D Revisiting Strategy 3 (Perturbation Consistency)
Researchers also observed that poisoned samples exhibit greater prediction consistency than those of benign samples under pixel-level amplification guo2023scale ; pal2024backdoor or weight-level alterations qi2023towards ; hou2024ibd . This implies that DNNs tend to overfit on triggers instead of learning semantic features.
Figure 6: The main pipeline of FLARE. Stage 1: Latent Representation Extraction: A backdoored model is trained on a poisoned dataset, and each training sample $x_{i}$ is forwarded to generate feature maps at hidden layers. The value ranges of each feature map are aligned using the statistics of the corresponding BN layers. An abnormal value is then extracted from each aligned feature map, and these values are aggregated across all hidden layers to form the final representation $\mathbf{R}_{i}$. Stage 2: poisoned samples Detection: With the representations $\mathbf{R}$ of all training samples, UMAP reduces $\mathbf{R}$ to $\hat{\mathbf{R}}$, followed by density-based clustering to separate samples into two clusters. The cluster with higher stability $\kappa^{C}$ is identified as poisoned.
@@ III-D Revisiting Strategy 3 (Perturbation Consistency)
Results. As shown in Figure 4, for BadNets (A2O), benign and poisoned samples exhibit significantly different confidence difference patterns, making them easily distinguishable. In contrast, under A2A and UT attacks, both benign and poisoned samples display similar confidence difference patterns (i.e., the confidence differences of poisoned samples also approach 1, similar to those of benign samples), indicating that perturbations significantly impact the predictions of poisoned samples. The sensitivity of poisoned samples to perturbations, similar to that of benign samples, indicates that DNNs struggle to overfit to poisoned samples in A2A and UT attacks. Thus, the assumption that backdoor connections are easier to learn than benign ones does not hold for A2A and UT attacks.
@@ III-E Revisiting Latent Separability on a Particular Layer
Previous studies chen2018detecting ; ma2022beatrix explored the latent separability between benign and poisoned samples in the feature space. In particular, these works primarily focused on the final hidden layer as a representative intermediate layer. We argue that their success also partly relied on the assumption that backdoor can be easily learned. In this section, we verify it.
Settings. We also conduct experiments using BadNets and WaNet on the CIFAR-10 dataset to facilitate our analysis. To assess latent separability, we use t-SNE visualization to examine the latent representations of both benign and poisoned samples across various hidden layers, with particular focus on shallow, middle, and deep layers. We select 2,000 random samples from the training set, including both benign and poisoned ones, with a poisoning rate of 0.1. All other settings are consistent with those described in Section III-C.
@@ IV-B2 Extracting Abnormal Features
As backdoor-related features usually function as dominant features, they tend to induce abnormally large or small activations in the feature map, as partly supported by findings in cai2022randomized . Building on this understanding, we propose to focus on identifying outliers within each feature map, specifically targeting the abnormally small or large values. In particular, all feature maps are normalized and their values adhere to a uniform distribution defined by the statistics of the BN layers after the previous alignment step. At this time, both abnormally large and small values fall outside the central distribution, corresponding to regions with low occurrence probability. This allows us to consistently extract the minimum values from each aligned feature map as indicators of backdoor-related features, without needing to separately select the largest or smallest values. Specifically, we can extract the minimum value (i.e., $r_{c}^{(l)}$) from the aligned output $\hat{a}^{(l)}$ of the $l$-th hidden layer via:
@@ IV-C1 Cluster Splitting
Here, $\mathcal{T}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$ denotes the dimensionality reduction function. In this paper, we choose the uniform manifold approximation and projection (UMAP) algorithm mcinnes2018umap as $\mathcal{T}$ for its ability to reduce high-dimensional data while preserving the topological structure of the original data manifold. FLARE then performs clustering on the reduced representation $\hat{\mathbf{R}}$: $\mathcal{C}=\mathcal{H}(\hat{\mathbf{R}})$, where $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$ represents the set of clusters identified by a clustering algorithm $\mathcal{H}$. In this paper, we use the hierarchical density-based spatial clustering of applications with noise (HDBSCAN) algorithm mcinnes2017hdbscan as $\mathcal{H}$ for its ability to handle clusters of varying shapes and densities. HDBSCAN constructs a hierarchy of clusters represented by a condensed tree $\mathbf{T}$, where clusters are partitioned across varying density levels $\lambda=1/d_{core}$, where $d_{core}$ denotes the reachability distance of an object to its ‘minPts-nearest’ neighbor. A higher density level corresponds to a smaller $d_{core}$, indicating that the samples in a cluster are closely packed together. By adjusting the density threshold, HDBSCAN can distinguish between dense poisoned samples and more widely distributed benign samples.
@@ IV-C2 Stable Subspace Selection
Given that poisoned samples typically share the same trigger-related features, they tend to aggregate into a tight cluster. In contrast, benign samples, originating from a more natural distribution, typically form compact clusters within each class while remaining separated across different classes. As a result, poisoned clusters remain stable across varying density levels, while benign clusters tend to fragment or shift as the density threshold increases. This difference makes cluster stability an effective indicator for identifying poisoned samples. Cluster stability measures the ability of a cluster to maintain its integrity as the density level $\lambda$ increases. With an increase in $\lambda$, the reachability distance $d_{core}$ between samples decreases, potentially causing distant samples to be assigned to different clusters. Traditionally, in HDBSCAN, cluster stability is calculated by aggregating the $\lambda$ of each sample within a cluster. However, in poisoned sample detection, where the number of poisoned samples is much smaller than that of benign samples, this approach introduces bias toward the majority class (benign samples). To alleviate this problem, we redefine cluster stability to focus on the persistence of a cluster across varying $\lambda$, as follows.
@@
To determine the optimal number of layers $k$ to exclude, FLARE designs an adaptive algorithm that dynamically selects a suitable $k$. The algorithm begins with a model configured with all $L$ hidden layers and progressively removes the last hidden layer of the modified model at each step. At each iteration, FLARE splits the condensed tree $\mathbf{T}$ at the root node to form two primary clusters and assesses the stability $\kappa^{c_{max}}$ of the larger cluster $c_{max}$. If the stability $\kappa^{c_{max}}$ surpasses a predefined threshold $\xi$, it suggests that the benign samples are cohesively grouped within the current subspace. To further ensure the cluster stability against potential anomalies, FALRE generates a condensed tree at each step and traverses from the root node of $c_{max}$ down to depth $d$. If the stability at any depth exceeds $\xi$, the current subspace is considered stable. The details of the selection process are provided in Algorithm 1.
@@ IV-C3 Stability-based Detection
Within this obtained subspace, FLARE splits the condensed tree $\mathbf{T}$ of the entire poisoned dataset at the root node to form two primary clusters, $C_{1}$ and $C_{2}$. FLARE then evaluates the cluster stability $\kappa^{C_{i}}$ for each cluster $C_{i}$ and identifies the cluster with the higher stability as poisoned. The set of poisoned samples $\mathcal{S}_{p}$ is:
@@
The detailed procedure for detecting poisoned training samples is presented in Algorithm 2.
Algorithm 1 Stable Subspace Selection
0:  backdoored model $\mathcal{F}$ with $L$ hidden layers; threshold $\xi$; maximum depth $d$.
0:  the stable representation subspace $\mathbf{R}^{\prime}$.
1:  for $k=0$ to $L-1$ do
2:     extract latent representations $\mathbf{R}\in\mathbb{R}^{N\times d}$ from $\mathcal{F}$;
3:     generate reduced representation $\mathbf{R}^{\prime}$ by excluding the last $k$ layers of $\mathbf{R}$, $\mathbf{R}^{\prime}=[\mathbf{r}^{1},\mathbf{r}^{2},\dots,\mathbf{r}^{L-k}]$;
4:     apply dimensionality reduction (e.g., UMAP) to get $\hat{\mathbf{R}}$ as in Eq. 6;
5:     apply clustering analysis (e.g., HDBSCAN) on $\hat{\mathbf{R}}$ to construct the condensed tree $\mathbf{T}$;
[…]
7:     while $Q$ is not empty do
8:        dequeue $(n,\,h)$ from $Q$;
9:        if $h\geq d$ then
10:           break
11:        end if
12:        $\lambda_{n}\leftarrow$ the density level of node $n$ in $\mathbf{T}$;
13:        if node $n$ has children then
14:           choose the child node $c_{\max}$ with the largest size;
15:           $\lambda_{c}\leftarrow$ the density level of node $c_{\max}$ in $\mathbf{T}$;
16:           compute the cluster stability $\kappa^{c_{max}}=\lambda_{c}-\lambda_{n}$;
17:           if $\kappa^{c_{max}}>\xi$ then
18:              return $\mathbf{R}^{\prime}$;
19:           end if
20:           enqueue $(c_{\max},\,h+1)$ into $Q$;
21:        end if
22:     end while
23:  end for
Algorithm 2 Stability-based Detection
0:  the stable representation subspace $\mathbf{R}^{\prime}$.
0:  the detected poisoned samples $\mathcal{S}_{p}$.
1:  apply dimensionality reduction (e.g., UMAP): $\hat{\mathbf{R}}=\mathcal{T}(\mathbf{R}^{\prime})$ as in Eq. 6;
2:  cluster $\hat{\mathbf{R}}$ to construct the condensed tree $\mathbf{T}$;
3:  split $\mathbf{T}$ at the root node into two clusters: $C_{1}$ and $C_{2}$;
4:  compute the stability $\kappa^{C_{i}}$ for each cluster $C_{i}$ according to Definition IV.1;
5:  identify the poisoned cluster $C_{p}$ with the higher stability score $\kappa^{C_{i}}$;
6:  return the detected poisoned samples $\mathcal{S}_{p}=\{x\mid x\in C_{p}\}$;
@@ V-A Main Settings
Datasets and Models. We conduct experiments on two classical image classification benchmark datasets, including CIFAR-10 krizhevsky2009learning and Tiny-ImageNet chrabaszcz2017downsampled . For the primary experiments, we employ ResNet-18 he2016deep as the default architecture. For ablation studies, we use VGG19 simonyan2014very and MobileNetV2 sandler2018mobilenetv2 .
Attacks Configurations. To validate the generalizability of our defense, we evaluate it against four representative types of backdoor attacks across three attack modes. These attacks are categorized as follows: (1) sample-agnostic attacks: BadNet gu2017badnets , Blend chen2017targeted , Trojan trojan2018 ; (2) sample-specific attacks: IAD nguyen2020input , WaNet nguyen2021wanet , ISSBA li2021invisible ; (3) clean-label attack: LC turner2019label ; (4) sparse attack: SIBA gao2024backdoor . The attack modes include the all-to-one (A2O), all-to-all (A2A), and untargeted (UT) paradigms. To comprehensively assess each attack mode, attacks are adapted accordingly. For example, BadNets is evaluated in three formats: BadNets (A2O), BadNets (A2A), and BadNets (UT). These attacks are implemented using the open-source BackdoorBox toolkit backdoorbox , strictly following the settings outlined in their papers. For all attacks, we set the target label $y_{t}$ to 0. Given the complexity of the Tiny-ImageNet dataset, we focus on the most representative attacks, omitting some similar-effect attacks to avoid excessive computational costs. To ensure consistently high attack success rates, we set the poisoning rate to 0.1 and the cover rate to 0.2.
TABLE I: The detection performance (%) of FLARE on the CIFAR-10 dataset using the ResNet-18 model. The best results (highest TPRs and lowest FPRs) are in bold, while TPRs $<$ 80% and FPRs $>$ 20% are marked in red as failed cases.
@@
TABLE II: The detection performance (%) of FLARE on the TinyImageNet dataset using the ResNet-18 model. The best results (highest TPRs and lowest FPRs) are in bold, while TPRs $<$ 80% and FPRs $>$ 20% are marked in red as failed cases.
@@
TABLE III: Effectiveness of secure training from scratch (FLARE (P)) on the “purified” CIFAR-10 dataset. The best results (highest BAs and lowest ASRs) are in bold, while all failed cases (BA drop or ASR $>$ 10%) are marked in red.
@@
TABLE IV: Effectiveness of secure training from scratch (FLARE (P)) on the “purified” Tiny-ImageNet dataset. The best results (highest BAs and lowest ASRs) are in bold, while all failed cases (BA drop or ASR $>$ 10%) are marked in red.
@@
Figure 7: Performance of our defense across different target labels of CIFAR-10. The consistently high TPR and low FPR across various target labels indicate stable and effective defense performance.
V-B Detection Performance
We mainly assess the effectiveness of FLARE against various backdoor attacks on the CIFAR-10 and Tiny-ImageNet datasets. The results, detailed in Table I and Table II, demonstrate that FLARE consistently achieves promising performance across different attack scenarios,
with the TPRs reaching 100% (or near 100%) while maintaining FPRs near 0%. The results demonstrate a significant improvement over baseline defenses, which fail in A2A and UT attack scenarios (highlighted in red). This failure is primarily due to their implicit assumptions that backdoor correlations are easier to learn than benign ones, which does not hold in more complex attack configurations. Additionally, FLARE maintains an FPR close to zero on datasets composed entirely of benign samples, indicating that it has minimal false positives.
V-C Effectiveness of Secure Training from Scratch
To assess the effectiveness of FLARE (P), which retrains models on a purified dataset with detected poisoned samples removed, we evaluated the BAs and ASRs of the retrained models on the CIFAR-10 and Tiny-ImageNet datasets. In particular, we define ‘failed cases’ as a BA drop of more than 10% compared to the original BA of the backdoored model without defense or an ASR greater than 10%, in line with the evaluation criteria used in existing defenses xu2024towards ; chen2025refine . As shown in Table III and Table IV, the ASRs of retrained models approach to nearly 0% in almost all cases. In contrast, all baseline methods fail in most cases, especially under all-to-all and untargeted attacks. In particular, we also observe some outliers with a high ASR, e.g., models under Blend (UT) and SIBA (UT) on CIFAR-10. This is mainly because even adding triggers of these attacks to a backdoor-free benign model may result in some ASR, instead of due to the failure of our method.
V-D Effectiveness of Backdoor Removal
To assess the effectiveness of FLARE (R), which directly removes backdoors by unlearning the detected poisoned samples, we also evaluated the BAs and ASRs of the purified models on the CIFAR-10 and Tiny-ImageNet datasets. As shown in Table VI and Table VII, FLARE (R) significantly reduces the ASRs of different attacks to nearly 0%, while the impact on BA remains minimal, generally around 1%. Compared to existing advanced defenses, FLARE (R) consistently proves effective across various attacks and datasets.
V-E Ablation Study
We conduct ablation studies to assess the impact of five specific factors on the effectiveness of FLARE: (1) target label for A2O attacks, (2) model architectures, (3) the module for selecting a stable subspace, (4) the hyper-parameters of the lambda threshold $\xi$ and maximum depth $d$, (5) the granularity in feature extraction.
TABLE V: Detection performance of FLARE on the CIFAR-10 dataset with different model architectures.
[…]
A2O BadNets 100.00 0.00 100.00 0.00
LC 100.00 0.00 100.00 0.00
ISSBA 99.98 0.00 99.90 0.00
A2A BadNets 100.00 0.00 100.00 0.00
UT BadNets 100.00 9.86 100.00 0.00
Impact of Model Architectures. In our main results, the model architecture is initially set to ResNet-18. To further assess the robustness of FLARE, we also evaluate it with VGG-19 and MobileNet v2 architectures. The results, shown in Table V, indicate that FLARE consistently achieves high TPRs, approaching 100%, while maintaining FPRs near 0%.
TABLE VI: Comparisons of FLARE (R) with SOTA backdoor-removal defenses on CIFAR-10. The best results (highest BAs and lowest ASRs) are in bold, while all failed cases (BA drop or ASR $>$ 10%) are marked in red.